On Tuesday, 20 of the 25 students in an algebra class took a test and their average was 86. On Wednesday, the other 5 students took the test, and their average was 81.What was the average for the entire class?
(A) 81.8 (B) 82 (C) 83.5 (D) 85 (E) 87

Answer by mananth(16949)

20 students average = 86
total of 20 students = 86 * 20 = 1720
..
5 students average = 81
total of 5 students = 5*81 = 405
..
Average of class = total of 25 / number
Average = 1720 +405 / 25
average = 2125/25
Average = 85 of the entire class
...
m.ananth@hotmail.ca